# Enhancing Image Layout Control with Loss-Guided Diffusion Models

### v1

@@ Abstract
Diffusion models are a powerful class of generative models capable of producing high-quality images from pure noise. In particular, conditional diffusion models allow one to specify the contents of the desired image using a simple text prompt. Conditioning on a text prompt alone, however, does not allow for fine-grained control over the composition and layout of the final image, which instead depends closely on the initial noise distribution. While most methods which introduce spatial constraints (e.g., bounding boxes) require fine-tuning, a smaller and more recent subset of these methods are training-free. They are applicable whenever the prompt influences the model through an attention mechanism, and generally fall into one of two categories. The first entails modifying the cross-attention maps of specific tokens directly to enhance the signal in certain regions of the image. The second works by defining a loss function over the cross-attention maps, and using the gradient of this loss to guide the latent. While previous work explores these as alternative strategies, we provide an interpretation for these methods which highlights their complimentary features, and demonstrate that it is possible to obtain superior performance when both methods are used in concert.
@@ 1 Introduction
An especially simple and intuitive way of describing a layout is to provide bounding boxes for various tokens in the text prompt. One way to realize a model taking such an input is to resort to training-based methods, wherein a pretrained model undergoes additional finetuning using training data where the images have been supplemented by their layouts. While such methods can achieve impressive performance Zheng et al. (2023); Zhang et al. (2023), this often involves the introduction of additional model complexity and training cost, in addition to the difficult task of compiling the training data. A recently proposed alternative approach Xie et al. (2023); Chen et al. (2024) uses the cross-attention module to achieve training-free layout control. By defining a loss over the attention maps, and by guiding or iteratively refining the latents to minimize this loss, it is possible to generate images that are reasonably faithful to the desired layout. One drawback, however, is that the ad-hoc choice of a loss function means that the image may be sampled from a low probability region of the learned distribution, particularly if the amount of guidance is high, as it competes with the denoising process described by the model’s learned score function. In this case, the sample suffers from poor image quality, even if the desired layout is achieved. It is also possible to directly modify the attention maps during the denoising process, to create images with high attention in the desired regions Balaji et al. (2023). This approach, like loss guidance, can also suffer from degraded image quality when the attention modification is sufficiently strong Chen et al. (2024).
In this paper, we propose injection loss guidance (iLGD), a framework in which the model’s denoising process maintains better agreement with loss guidance, such that sampled images simultaneously present the appropriate layout and maintain good image quality. First, we bias the latent towards the desired layout by altering the model’s attention maps directly, a process which we refer to as attention injection. The goal of injection is not to produce a latent which perfectly adheres to the desired layout, but rather, to produce a latent which is close enough to the desired layout that we can afford to use a smaller amount of loss guidance. In doing so, we reimagine the role of injection as a coarse biasing of the diffusion process, and loss guidance as a refiner. We show that such a framework is capable of controlling the image layout, while achieving superior image quality in many cases.
@@ 2.1 Generative Models
Generative models learn to estimate a data distribution with the goal of generating samples from this distribution. Autoregressive models Oord et al. (2016); Van Den Oord et al. (2016); Salimans et al. (2017); Germain et al. (2015) are likelihood-based models that predict a distribution over each pixel in an ordered sequence, given some window of previous pixels. While they are capable of modelling complex dependencies between different variables, they can be computationally expensive to train and inefficient to sample from, particularly when dealing with longer sequences. Generative adversarial networks (GANs) Goodfellow et al. (2014) are another type of generative model that are based on adversarial training, and are capable of producing high quality samples. However, they are often prone to mode collapse and are sometimes difficult to train. Variational Autoencoders (VAEs) Kingma and Welling (2014) are easier to train, but are known to produce blurry samples. Flow models Rezende and Mohamed (2015) similarly suffer from lower quality samples, although, like GANs and VAEs, they are efficient to sample from.
More recently, a new family of generative models, known as diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song and Ermon (2019), have achieved superior results on image synthesis compared to the previous state of the art Dhariwal and Nichol (2021), with samples exhibiting incredible diversity and image quality. While early diffusion models were formulated as Markov chains and relied on a large number of transitions to generate samples, Song et al. (2021) showed that such a sampling procedure can be viewed as a discretization of a certain stochastic differential equation (SDE). In particular, Song et al. (2021) showed that there exists a family of SDEs, whose solutions are sampling trajectories from the diffusion model. One such SDE, called the probability flow ODE, is completely deterministic and contains no noise. This enables the use of various ordinary differential equation (ODE) solvers for efficient sampling Karras et al. (2022); Lu et al. (2022).
@@ 2.2 Controllable Generation
Diffusion models can use a wide range of techniques for controllable generation. SDEdit Meng et al. (2022) allows the user to specify a layout by using paint strokes, which are noised to time $t<T$ to provide an initialization for solving the reverse-time SDE. The realism of the final image is, however, sensitive to the initial noise level $\sigma_{t}$, and guiding the generation of new images requires all high-level features to be specified in the stroke image. A more user-friendly method by  Voynov et al. (2023) requires only a simple sketch to guide the denoising process. The user-provided sketch is compared with the edges extracted by a latent edge predictor in order to compute a loss, which is used to iteratively refine the latent. In this case, the latent edge predictor must be trained, and sketches of more complicated scenes may be tedious. Zhang et al. (2023) propose a more general method in which a separate encoder network takes as input a conditioning control image, such as a sketch, depth map, or scribble, to guide the generation process. Unfortunately, this requires finetuning a large pretrained encoder. When the inputs are specified as bounding boxes, smaller trainable modules can be used between the layers of the denoising UNet to encode layout information Zheng et al. (2023). Cheng et al. (2023) also use an additional module which takes in bounding box inputs, injecting it directly after the self-attention layers. Once again, however, neither method can be adapted immediately, as the additional parameters necessitate further training.
@@
A number of other methods have been proposed that are training-free. Bansal et al. (2024) define a generic loss on the noiseless latent $\hat{\mathbf{z}}_{0}$ predicted from $\mathbf{z}_{t}$ by Tweedie’s formula, and subsequently perform loss guidance on the latent $\mathbf{z}_{t}$ at each time step. One downside of this method is that, while the loss expects clean images, the predicted latent $\hat{\mathbf{z}}_{0}$ is an approximation to only an average of possible generated images, and so can be blurry for large times $t$. An alternative approach was proposed by Bar-Tal et al. (2023), who used separate score functions on various regions in a latent diffusion model, with an optimization step at each iteration designed to fuse the separate diffusion paths. While this method can be effective in many cases, it can nonetheless exhibit patchwork artifacts, where the final image appears to be composed of several images rather than depicting a single scene.
@@
Alternatively, several works have explored the use of cross-attention to achieve training-free layout control. Hertz et al. (2023) demonstrate how attention maps can be injected from one diffusion process to another, and reweighted to control the influence of specific tokens. Subsequent work by Balaji et al. (2023) builds upon this idea by directly manipulating the values in the attention map to obtain the desired layout, although it is difficult to precisely localize objects appearing in an image with this method alone.  Singh et al. (2023) also use this technique to improve semantic control in stroke-guided image synthesis, which they combine with loss guidance based on the stroke image to improve the realism of generated images. Instead of using strokes, Chen et al. (2024) show that controlling layout with bounding boxes is possible by using loss guidance, where the loss is defined on the attention maps, although this method requires searching for a suitable noise initialization. Concurrent work by Xie et al. (2023) and Couairon et al. (2023) also use attention-based loss guidance; the former adds spatial constraints to control the scale of the generated content, while the latter uses segmentation maps instead of bounding boxes. Epstein et al. (2023) show that it is even possible to control properties of objects in an image, such as their shape, size, and appearance, through their attention maps, and subsequently manipulate these properties through loss guidance.
@@ 3 Preliminaries
3.1 Diffusion Models
Diffusion models Ho et al. (2020) are characterized by two principle algorithms. The first is the forward process, wherein the data $\mathbf{x}_{0}$ is gradually corrupted by Gaussian noise until it becomes pure noise, which we denote by $\mathbf{x}_{T}$. The reverse process moves in the opposite direction, attempting to recover the data by iteratively removing noise. The denoiser $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is typically a UNet Ronneberger et al. (2015) which accepts an image $\mathbf{x}_{t}$, and predicts its normalized noise content $\bm{\epsilon}$. Removing a fraction of this noise yields a slightly denoised image $\mathbf{x}_{t-1}$. Repeating this process over $T$ steps produces a noise-free image $\mathbf{x}_{0}$.
Operating directly on the image $\mathbf{x}_{t}$ in pixel-space is computationally expensive. As an alternative, latent diffusion models have been proposed to curtail this high cost, in which the denoising procedure is performed in latent space, whose dimensionality is typically much lower than pixel space. Stable Diffusion Rombach et al. (2022) is one example of a latent diffusion model which achieves state-of-the-art performance on various image synthesis tasks. It leverages a powerful autoencoder to project to and from latent space, where the standard denoising procedure is performed. Images in latent space are typically denoted by $\mathbf{z}_{t}$, and the encoder and decoder are denoted by $\mathcal{E}$ and $\mathcal{D}$, respectively, so that $\mathbf{z}_{t}=\mathcal{E}(\mathbf{x}_{t})$ and $\mathbf{x}_{t}=\mathcal{D}(\mathbf{z}_{t})$.
During training, samples from the true data distribution $q(\mathbf{x}_{0})$ are corrupted via the forward process. By training a diffusion model to learn a reverse process in which it iteratively reconstructs these noisy samples into noise-free samples, it is possible to generate images from pure noise at inference time. This corresponds to sampling from an approximation $p_{\theta}(\mathbf{x}_{0})$ to the data distribution, $q(\mathbf{x}_{0})$. This generation process can be guided by introducing an additional input vector $\mathbf{y}$, which is often a text prompt. In this case, the model produces samples from an approximation $p_{\theta}(\mathbf{x}_{0}|\mathbf{y})$ to the conditional distribution $q(\mathbf{x}_{0}|\mathbf{y})$. A more detailed discussion of diffusion models is provided in Appendix A.
@@ 3.2 Cross-Attention
$\mathbf{X}\mathbf{W_{q}}=\mathbf{Q}\in\mathbbm{R}^{n\times d_{k}},$ (1)
$\mathbf{Y}\mathbf{W_{k}}=\mathbf{K}\in\mathbbm{R}^{k\times d_{k}},$ (2)
$\mathbf{Y}\mathbf{W_{v}}=\mathbf{V}\in\mathbbm{R}^{k\times d_{v}}.$ (3)
Subsequently, the attention weights are computed as
$A=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d_{k}}}\right)\in\mathbbm{R}^{n\times k}.$ (4)
@@
$\mathbf{Z}=A\mathbf{V}\in\mathbbm{R}^{n\times d_{v}}.$ (5)
@@ 3.3 Score Matching
The forward and reverse processes can be modelled by solutions of stochastic differential equations (SDEs) Song et al. (2021). While determining the coefficients of the forward process SDE is straightforward, the reverse process corresponds to a solution of the reverse-time SDE, which requires learning the score of the intractable marginal distribution $q(\mathbf{x}_{t})$. Instead, Song et al. (2021) use the score-matching objective
@@
for some nonnegative function $\lambda\colon[0,T]\to\mathbbm{R}$.
@@
where $\sigma_{t}$ is the standard deviation of the forward process at time $t$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ predicts the normalized noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in $\mathbf{x}_{t}$.
@@
3.4 Classifier-Free Guidance
In order to generate images following a user-supplied text prompt, the denoiser $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ of a latent diffusion model is trained with an additional input given by a sequence of token embeddings $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{k}\}$. A single denoiser, usually a UNet, is trained over a variety of text prompts, and the token embeddings influence the denoiser by a cross-attention mechanism in both the contractive and expansive layers. Ho and Salimans (2021) found that, rather than sampling images using the conditional denoiser alone, better results can be obtained by taking a combination of conditional and unconditional noise estimates,
In this paper, we denote the CFG noise prediction by $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$, or sometimes just $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$, when the dependence on $\mathbf{y}$ is clear.
@@ 3.5 Controllable Layout Generation
We use BoxDiff Xie et al. (2023) as our primary point of comparison. In their paper, the authors apply spatial constraints on the attention maps of a latent diffusion model to derive a loss, and directly update the latent at time step $t$ by replacing it with
$\mathbf{z}^{\prime}_{t}=\mathbf{z}_{t}-\alpha_{t}\cdot\nabla_{\mathbf{z}_{t}}\mathcal{L}.$ (9)
The parameter $\alpha_{t}$ decays linearly with $t$, and controls the strength of the loss guidance in each iteration. The loss $\mathcal{L}$ is computed as the sum of three separate losses, $\mathcal{L}=\mathcal{L}_{\text{IB}}+\mathcal{L}_{\text{OB}}+\mathcal{L}_{\text{CC}}$, where $\mathcal{L}_{\text{IB}}$ encourages the attentions of the target tokens to be large inside the bounding boxes, $\mathcal{L}_{\text{OB}}$ penalizes attentions appearing outside of the bounding boxes, and $\mathcal{L}_{\text{CC}}$ encourages higher attentions near the corners of the bounding boxes.
@@ 4.1 Attention Injection
Hertz et al. (2023) observed that, by extracting the attention maps from a latent diffusion process and applying them to another one with a modified token sequence, it is possible to transfer the composition of an image. This technique is very effective, but requires an original set of attention maps which produce the desired layout, and it is not feasible to generate images until this layout is obtained.
Instead, we rely on the observation that the attention maps early in the diffusion process are strong indicators of the generated image’s composition. For early timesteps, both the latents and the attention maps are relatively diffuse, and don’t suggest any fine details about objects in the image. Motivated by this, we manipulate the attention maps by artificially enhancing the signal in certain regions. Given a list of target tokens $S\subset\{1,2,\ldots,k\}$, we define $m=\{\mathbf{m}_{1},\mathbf{m}_{2},\ldots,\mathbf{m}_{k}\}\in\mathbbm{R}^{n\times k}$ by setting each mask $\mathbf{m}_{j}$, $j\in S$, equal to $1$ over the region that the text token $\mathbf{y}_{j}$ should correspond to, and zero otherwise, and perform injection by replacing the cross-attention map $A_{t}$ at time $t$ with
$A^{\prime}_{t}=\text{softmax}\left(\frac{\mathbf{Q}_{t}\mathbf{K}_{t}^{T}+\nu_{t}m}{\sqrt{d_{k}}}\right),$ (10)
for some $\nu_{t}>0$. We follow Balaji et al. (2023) and use the scaling

### v2

@@ Abstract
Diffusion models are a powerful class of generative models capable of producing high-quality images from pure noise using a simple text prompt. While most methods which introduce additional spatial constraints into the generated images (e.g., bounding boxes) require fine-tuning, a smaller and more recent subset of these methods take advantage of the models’ attention mechanism, and are training-free. These methods generally fall into one of two categories. The first entails modifying the cross-attention maps of specific tokens directly to enhance the signal in certain regions of the image. The second works by defining a loss function over the cross-attention maps, and using the gradient of this loss to guide the latent. While previous work explores these as alternative strategies, we provide an interpretation for these methods which highlights their complimentary features, and demonstrate that it is possible to obtain superior performance when both methods are used in concert.
@@ 1 Introduction
An especially simple and intuitive way of describing a layout is to provide bounding boxes for various tokens in the text prompt. One way to realize a model taking such an input is to resort to training-based methods, wherein a pretrained model undergoes additional finetuning using training data where the images have been supplemented by their layouts. While such methods can achieve impressive performance Zheng et al. (2023); Zhang et al. (2023), this often involves the introduction of additional model complexity and training cost, in addition to the difficult task of compiling the training data. A recently proposed alternative approach Xie et al. (2023); Chen et al. (2024) uses the cross-attention module to achieve training-free layout control.
In this paper, we propose injection loss guidance (iLGD), a training-free framework in which the model’s denoising process works in synergy with loss guidance, such that sampled images simultaneously present the appropriate layout and maintain good image quality. First, we bias the latent towards the desired layout by altering the model’s attention maps directly, a process which we refer to as attention injection. The goal of injection is not to produce a latent which perfectly adheres to the desired layout, but rather to provide a comfortable starting point for loss guidance. The resulting latent is sufficiently close to the desired layout that we can afford to use a smaller amount of loss guidance. In doing so, we reimagine the role of injection as a coarse biasing of the diffusion process, and loss guidance as a refiner. We show that such a framework is capable of controlling the layout, while achieving superior image quality in many cases.
@@ 2.1 Generative Models
Generative models learn to estimate a data distribution with the goal of generating samples from this distribution. Recently , a new family of generative models, known as diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song and Ermon (2019), have achieved superior results on image synthesis compared to the previous state of the art Dhariwal and Nichol (2021), with samples exhibiting incredible diversity and image quality. While early diffusion models were formulated as Markov chains and relied on a large number of transitions to generate samples, Song et al. Song et al. (2021) showed that such a sampling procedure can be viewed as a discretization of a certain stochastic differential equation (SDE). In particular, Song et al. Song et al. (2021) showed that there exists a family of SDEs, whose solutions are sampling trajectories from the diffusion model. One such SDE, called the probability flow ODE, is completely deterministic and contains no noise. This enables the use of various ordinary differential equation (ODE) solvers for efficient sampling Karras et al. (2022); Lu et al. (2022).
@@ 2.2 Controllable Generation
Diffusion models can use a wide range of techniques for controllable generation. SDEdit Meng et al. (2022) allows the user to specify a layout by using paint strokes, which are noised to time $t<T$ to provide an initialization for solving the reverse-time SDE. The realism of the final image is, however, sensitive to the initial noise level $\sigma_{t}$, and guiding the generation of new images requires all high-level features to be specified in the stroke image. A more user-friendly method by Voynov et al. Voynov et al. (2023) requires only a simple sketch to guide the denoising process. The user-provided sketch is compared with the edges extracted by a latent edge predictor in order to compute a loss, which is used to iteratively refine the latent. In this case, the latent edge predictor must be trained, and sketches of more complicated scenes may be tedious. Zhang et al. Zhang et al. (2023) propose a more general method in which a separate encoder network takes as input a conditioning control image, such as a sketch, depth map, or scribble, to guide the generation process. Unfortunately, this requires finetuning a large pretrained encoder. When the inputs are specified as bounding boxes, smaller trainable modules can be used between the layers of the denoising UNet to encode layout information Zheng et al. (2023). Cheng et al. Cheng et al. (2023) also use an additional module which takes in bounding box inputs, injecting it directly after the self-attention layers. Once again, however, neither method can be adapted immediately, as the additional parameters necessitate further training.
@@
A number of other methods have been proposed that are training-free. Bansal et al. Bansal et al. (2024) define a generic loss on the noiseless latent $\hat{\mathbf{z}}_{0}$ predicted from $\mathbf{z}_{t}$ by Tweedie’s formula, and subsequently perform loss guidance on the latent $\mathbf{z}_{t}$ at each time step. One downside of this method is that, while the loss expects clean images, the predicted latent $\hat{\mathbf{z}}_{0}$ is an approximation to only an average of possible generated images, and so can be blurry for large times $t$. An alternative approach called MultiDiffusion was proposed by Bar-Tal et al. Bar-Tal et al. (2023), who used separate score functions on various regions in a latent diffusion model, with an optimization step at each iteration designed to fuse the separate diffusion paths. While this method can be effective in many cases, it can nonetheless exhibit patchwork artifacts, where the final image appears to be composed of several images rather than depicting a single scene.
@@
Alternatively, several works have explored the use of cross-attention to achieve training-free layout control. Hertz et al. Hertz et al. (2023) demonstrate how attention maps can be injected from one diffusion process to another, and reweighted to control the influence of specific tokens. Subsequent work by Balaji et al. Balaji et al. (2023) builds upon this idea by directly manipulating the values in the attention map to obtain the desired layout, although it is difficult to precisely localize objects appearing in an image with this method alone. Singh et al. Singh et al. (2023) also use this technique to improve semantic control in stroke-guided image synthesis, which they combine with loss guidance based on the stroke image to improve the realism of generated images. Instead of using strokes, Chen et al. Chen et al. (2024) show that controlling layout with bounding boxes is possible by using loss guidance, where the loss is defined on the attention maps, although this method requires searching for a suitable noise initialization. Concurrent work by Xie et al. Xie et al. (2023) and Couairon et al. Couairon et al. (2023) also use attention-based loss guidance; the former adds spatial constraints to control the scale of the generated content, while the latter uses segmentation maps instead of bounding boxes. Epstein et al. Epstein et al. (2023) show that it is even possible to control properties of objects in an image, such as their shape, size, and appearance, through their attention maps, and subsequently manipulate these properties through loss guidance.
@@ 3.1 Cross-Attention
$\mathbf{X}\mathbf{W_{q}}=\mathbf{Q}\in\mathbbm{R}^{n\times d_{k}},$
$\mathbf{Y}\mathbf{W_{k}}=\mathbf{K}\in\mathbbm{R}^{k\times d_{k}},$
$\mathbf{Y}\mathbf{W_{v}}=\mathbf{V}\in\mathbbm{R}^{k\times d_{v}},$
where $d_{k}$ denotes the dimension of the query and key vectors, and $d_{v}$ denotes the dimension of the value vectors. Subsequently, the attention weights are computed as
$A=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d_{k}}}\right)\in\mathbbm{R}^{n\times k}.$
@@
$\mathbf{Z}=A\mathbf{V}\in\mathbbm{R}^{n\times d_{v}}.$
@@ 3.2 Score Matching
The forward and reverse processes can be modelled by solutions of stochastic differential equations (SDEs) Song et al. (2021). While determining the coefficients of the forward process SDE is straightforward, the reverse process corresponds to a solution of the reverse-time SDE, which requires learning the score of the intractable marginal distribution $q(\mathbf{x}_{t})$. Instead, Song et al. Song et al. (2021) use the score-matching objective
@@
for some positive function $\lambda\colon[0,T]\to\mathbbm{R}$.
@@
where $\sigma_{t}$ is the standard deviation of the forward process at time $t$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ predicts the scaled noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in $\mathbf{x}_{t}$.
@@
In practice, the denoising process usually operates on a latent variable $\mathbf{z}_{t}$, where an autoencoder is used to project to and from image space $\mathbf{x}_{t}$. In order to generate images which correspond to a user-supplied text prompt, the noise predictor $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ is conditioned on text prompts $\mathbf{y}$, and samples are computed using this noise predictor by classifier-free guidance (CFG). In this paper, we denote the CFG noise prediction by $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$, or sometimes just $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$, when the dependence on $\mathbf{y}$ is clear. We provide more details about diffusion models in Appendix A.
@@ 3.3 Controllable Layout Generation
We use BoxDiff Xie et al. (2023), the method of Chen et al. Chen et al. (2024), and MultiDiffusion Bar-Tal et al. (2023) as our primary points of comparison. In Xie et al. (2023) and Chen et al. (2024), the authors apply spatial constraints on the attention maps of a latent diffusion model to derive a loss, and directly update the latent at time step $t$ by replacing it with
$\mathbf{z}^{\prime}_{t}=\mathbf{z}_{t}-\alpha_{t}\cdot\nabla_{\mathbf{z}_{t}}\mathcal{L}_{\mathbf{y}}(\mathbf{z}_{t}),$ (3)
where $\mathcal{L}_{\mathbf{y}}$ is a loss function which depends on the neural network, the prompt $\mathbf{y}$, and a set of bounding boxes. The parameter $\alpha_{t}$ controls the strength of the loss guidance in each iteration. In BoxDiff, $\alpha_{t}$ decays linearly from $t=T$ to $t=0$, while in Chen et al., $\alpha_{t}=\eta\cdot\sigma_{t}^{2}$ for some scale factor $\eta>0$. In MultiDiffusion, letting $\Phi_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ denote the function which computes $\mathbf{z}_{t-1}$ from $\mathbf{z}_{t}$ using the predicted scaled noise ${\bm{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$, a new latent $\mathbf{z}^{\prime}_{t-1}$ is produced by solving the optimization problem
$\mathbf{z}^{\prime}_{t-1}=\mathop{\mathrm{arg\,min}}_{\mathbf{z}}\sum_{i=1}^{k}\bigl{\|}\mathbf{m}_{i}\otimes\bigl{(}\mathbf{z}-\Phi_{\theta}(\mathbf{z}_{t},t,\mathbf{y}_{i})\bigr{)}\bigr{\|}^{2},$
where each $\mathbf{m}_{i}\in\{0,1\}^{n}$ is a mask corresponding to the bounding box of the prompt $\mathbf{y}_{i}$.
@@ 4.1 Attention Injection
Hertz et al. Hertz et al. (2023) observed that, by extracting the attention maps from a latent diffusion process and applying them to another one with a modified token sequence, it is possible to transfer the composition of an image. This technique is very effective, but requires an original set of attention maps which produce the desired layout, and it is not feasible to generate images until this layout is obtained.
Instead, we rely on the observation that the attention maps early in the diffusion process are strong indicators of the generated image’s composition. For early timesteps, both the latents and the attention maps are relatively diffuse, and don’t suggest any fine details about objects in the image. Motivated by this, we manipulate the attention maps by artificially enhancing the signal in certain regions. Given a list of target tokens $S\subset\{1,2,\ldots,k\}$, we define $\mathbf{m}=\{\mathbf{m}_{1},\mathbf{m}_{2},\ldots,\mathbf{m}_{k}\}\in\mathbbm{R}^{n\times k}$ by setting each mask $\mathbf{m}_{j}$, $j\in S$, equal to $1$ over the region that the text token $\mathbf{y}_{j}$ should correspond to, and zero otherwise, and perform injection by replacing the cross-attention map $A_{t}$ at time $t$ with
$A^{\prime}_{t}=\text{softmax}\left(\frac{\mathbf{Q}_{t}\mathbf{K}_{t}^{T}+\nu_{t}\mathbf{m}}{\sqrt{d_{k}}}\right),$ (4)
where $\nu_{t}>0$. We follow Balaji et al. Balaji et al. (2023) and use the scaling